# arXiv:2603.00917v1[cs.CL] 01 Mar 2026

### v1

@@ 4.2 Effect of Prompt Style on Accuracy
Prompt style demonstrated a consistent and measurable effect on model accuracy across all five models and three datasets (Figure 3). The Roleplay prompt style — in which models were instructed to respond as a practicing physician — consistently underperformed relative to all other styles. The most pronounced decline was observed in Phi-3 Mini, which achieved 48.0% accuracy under the Direct style but only 26.5% under Roleplay on MedQA, representing a drop of 21.5 percentage points. Similar reductions were observed across Llama 3.2, Gemma 2, and Mistral 7B, confirming that roleplay prompting is systematically detrimental rather than model-specific (Figure 4).
In contrast, the Direct and Original prompt styles produced the highest and most stable accuracy across models and datasets, suggesting that minimal framing is preferable for reliable clinical question answering. These findings carry direct implications for healthcare AI deployment: prompt formulations intended to simulate clinical expertise — such as persona-based instructions — may paradoxically reduce the reliability of small language model outputs in patient-facing applications.
@@ 4.3 Instruction-Following Failure Rate
The UNKNOWN rate — the proportion of responses that did not contain a valid answer option — varied substantially across models and datasets (Figure 5). Phi-3 Mini exhibited the highest UNKNOWN rate on MedQA at 10.5%, indicating that approximately one in ten queries failed to produce a usable response. Mistral 7B, despite being the largest model evaluated, also demonstrated non-negligible UNKNOWN rates across all three datasets (4.7%–7.2%), suggesting that model size alone does not guarantee reliable instruction following.
In contrast, Llama 3.2 and Gemma 2 achieved the lowest UNKNOWN rates across datasets (0.8%–2.8% and 0.9%–2.1% respectively), consistently producing responses in the expected format. These results suggest that instruction adherence is a model-specific characteristic independent of parameter count.
Meditron-7B exhibited the most severe instruction-following failures of all models evaluated, with UNKNOWN rates of 22.8% on MedQA, 22.4% on MedMCQA, and 99.0% on PubMedQA. The near-complete failure on PubMedQA — where 99 of 100 responses were invalid — is consistent with Meditron’s architecture as a base pretraining model without instruction tuning. Unlike the instruction-following failures observed in Phi-3 Mini, which appear to be prompt-style dependent, Meditron’s failures reflect a fundamental absence of structured response capability rather than prompt sensitivity.
@@ 4.4 Consistency Versus Accuracy Relationship
Analysis of the relationship between mean consistency scores and overall accuracy revealed no clear positive correlation across models and datasets (Figure 6). Gemma 2 exemplifies this disconnect most starkly: it achieved the highest consistency scores across all datasets (0.845–0.888) while simultaneously recording the lowest accuracy (33.0–43.5%), indicating that it consistently produces incorrect answers with high confidence. Conversely, Llama 3.2 demonstrated moderate consistency (0.774–0.807) yet achieved the highest accuracy across most datasets (49.0–65.0%), showing that a model can be more frequently correct even when its answers vary somewhat across prompt styles.
The distribution of consistency scores further illustrates model-level differences in reliability (Figure 7). Phi-3 Mini exhibited the widest spread of consistency scores on MedQA, indicating high variability in prompt sensitivity across questions. Gemma 2 showed a tighter, higher distribution — consistently stable but systematically inaccurate.
@@ 5.1 Consistency and Accuracy are Independent Metrics
A central finding of this study is that consistency and accuracy represent independent dimensions of small language model performance in clinical question answering. While prior evaluations of clinical AI have predominantly focused on accuracy as the primary metric (Bedi and others, 2025), our results demonstrate that a model can achieve high consistency while remaining systematically inaccurate — a failure mode we term reliable incorrectness.
Gemma 2 illustrates this most clearly, achieving the highest mean consistency scores across all datasets (0.845–0.888) while recording the lowest overall accuracy (33.0–43.5%). This pattern suggests that the model has learned stable response tendencies that are nonetheless clinically wrong.
The danger of reliable incorrectness is as much psychological as technical. Consider a primary care physician using an AI assistant to support diagnostic reasoning. If the model returns the same answer regardless of how a question is phrased, this repetition creates an illusion of reliability. A clinician might reason: “it keeps suggesting pulmonary embolism — it must be confident.” That apparent stability feels reassuring. But if the correct diagnosis is pneumonia, the model’s consistency actively reinforces a false conclusion. Over time, clinicians may develop misplaced trust in systems that appear stable and decisive without recognising that the stability reflects systematic error rather than clinical knowledge. In practice, this could translate into repeated misdiagnoses, inappropriate investigations, or delayed treatment — not because the model behaves randomly, but because it is confidently and predictably incorrect.
This finding aligns with Kim et al. (Kim and others, 2025), who demonstrate that foundation models frequently produce hallucinations with high apparent confidence, and that clinicians consider such outputs particularly dangerous precisely because they are difficult to detect. Our results extend this concern to the domain of prompt sensitivity: a model that produces consistent but wrong answers across varied prompt formulations compounds this risk by appearing robust to rephrasing.
These observations underscore that accuracy and consistency must be evaluated jointly in any clinical AI assessment framework. Neither metric alone is sufficient to characterise the safety profile of a model intended for healthcare deployment.
@@ 5.2 Roleplay Prompts Consistently Underperform
We propose that roleplay prompts introduce a form of task interference that is particularly detrimental for small language models. When a prompt begins with persona framing such as “You are a senior physician taking a licensing examination,” the model must simultaneously interpret and simulate a clinical identity while reasoning through a structured medical question. For smaller models with limited representational capacity, this dual requirement may dilute attention from the core reasoning task. Furthermore, persona-based prompts may activate patterns learned from conversational or narrative training data rather than the structured exam-style reasoning required for clinical QA benchmarks — effectively shifting the task from question answering to performance-based text generation.
This finding has direct practical implications. Prompt engineering guidelines for clinical AI systems often recommend persona framing to make outputs sound more authoritative or clinically appropriate (White and others, 2023). Our results suggest that for small open-source models, such framing may paradoxically reduce factual reliability. Developers building clinical decision support tools on small language models should favour minimal, direct prompt formulations over persona-based instructions.
This observation also extends findings from Ngweta et al. (Ngweta and others, 2024), who demonstrated that prompt format changes cause measurable performance variation in general NLP tasks. Our results show that in clinical settings, one specific format — roleplay — is systematically and substantially more harmful than others, representing a concrete and actionable finding for healthcare AI deployment.
@@ 5.3 Implications for Low-Resource Healthcare Deployment
A primary motivation for this study was the practical challenge of deploying AI in resource-constrained healthcare environments. Many rural clinics, community hospitals, and health systems in low-income settings lack access to GPU infrastructure, reliable high-speed internet, or cloud-based AI services. In these contexts, locally deployable open-source models in the 2B–7B parameter range represent a more realistic and equitable alternative to large proprietary systems. Our experimental design intentionally reflected these constraints: all models were evaluated on consumer CPU hardware without domain-specific fine-tuning, simulating realistic deployment conditions rather than idealised research environments.
Our findings provide concrete guidance for model selection in such settings. Although Gemma 2 demonstrated the highest consistency scores, its substantially lower accuracy suggests it may systematically mislead clinicians — a risk that is amplified in low-oversight environments where AI outputs may not be routinely verified. Mistral 7B, despite being the largest model evaluated, did not clearly outperform smaller alternatives and exhibited unexpected instruction-following failures, suggesting that larger parameter counts do not automatically justify greater hardware demands.
Llama 3.2 (3B) emerged as the most balanced candidate across our evaluation criteria, achieving the highest overall accuracy (49.0–65.0%), moderate consistency, and among the lowest instruction-following failure rates. For low-resource healthcare deployment scenarios where clinician oversight may be limited, a model offering strong accuracy and reliable instruction adherence is preferable to one that maximises consistency at the expense of correctness.
More broadly, our results suggest that deployment decisions for clinical AI systems should be guided by joint evaluation of accuracy, consistency, and instruction compliance rather than any single metric. In settings where errors carry direct patient safety implications, the cost of systematic incorrectness — even when delivered consistently — outweighs the benefit of apparent stability. This study provides a practical evaluation framework for low-resource healthcare AI deployment that can be reproduced without specialised infrastructure, making it accessible to researchers and practitioners in exactly the settings it is designed to serve.
@@ 5.4 Domain Knowledge Versus Instruction Following
This finding suggests that domain knowledge and task usability are orthogonal capabilities. A model can encode rich medical knowledge through continued pretraining yet remain clinically unusable for structured question answering without the ability to follow task-specific instructions. For practitioners deploying AI in clinical settings, this distinction has direct practical implications: selecting a model based on its training data alone — without verifying instruction-following capability — may result in a system that is knowledgeable but functionally unusable in structured workflows.
This also motivates a clear direction for future work: instruction-tuning Meditron or similar domain-pretrained models on structured clinical QA data, and evaluating whether the combination of domain knowledge and instruction following produces models that are both more accurate and more consistent than general instruction-tuned alternatives.
@@ 5.5 Limitations
Several limitations should be considered when interpreting these findings. First, we sampled 200 questions per dataset, which provides a computationally feasible and reproducible evaluation but may not fully capture the diversity of each benchmark. Second, the three datasets used — MedQA, MedMCQA, and PubMedQA — focus primarily on structured examination-style and research-based questions, which may not reflect the complexity of real-world clinical dialogue or unstructured patient interactions.
Third, we did not conduct human clinical evaluation to assess whether consistent but incorrect model outputs would meaningfully influence physician decision-making in practice. Such evaluation would strengthen the clinical validity of our findings but was beyond the scope of this study. Fourth, all models were evaluated without domain-specific fine-tuning, which may underestimate performance achievable through medical adaptation techniques such as instruction tuning or retrieval-augmented generation (Lewis and others, 2020).
Fifth, our evaluation was restricted to multiple-choice and yes/no/maybe answer formats, whereas real clinical environments frequently involve open-ended reasoning, differential diagnosis, and contextual patient data. Sixth, hardware constraints limited our evaluation to models in the 2B–7B parameter range running on consumer CPU hardware, precluding direct comparison with larger proprietary systems such as GPT-4 or Claude. Seventh, while our five prompt styles were designed to simulate realistic variation in clinical query formulation, they do not exhaust the full range of phrasings a clinician might use in practice. Finally, Meditron-7B was evaluated in its base pretraining form without instruction tuning, which limits direct comparison with instruction-tuned models and may underrepresent the potential of domain-adapted models when properly instruction-tuned for structured clinical tasks.
Together, these limitations suggest that our results should be interpreted as an initial systematic analysis of prompt sensitivity under constrained deployment conditions rather than a comprehensive evaluation of all clinical AI scenarios. Future work should address these gaps through larger samples, broader dataset coverage, and human clinical validation studies.
@@ 5.6 Future Work
Several directions emerge naturally from this study. First, future research should evaluate whether domain-specific fine-tuning of small models on medical corpora reduces prompt sensitivity and improves consistency. If fine-tuned models demonstrate stronger stability under prompt variation, this would suggest that training data alignment — rather than model architecture — is the primary driver of inconsistency.
In particular, instruction-tuning Meditron-7B or similar domain-pretrained models on structured clinical QA data represents a direct next step — testing whether combining domain knowledge with instruction-following capability produces models that are both more accurate and more consistent than general instruction-tuned alternatives evaluated in this study.
Second, integrating retrieval-augmented generation (RAG) pipelines with small clinical models represents a promising direction. RAG-augmented systems ground responses in retrieved evidence, which may reduce both factual hallucinations and prompt-induced answer variation. Testing whether external knowledge retrieval mitigates the consistency-accuracy tradeoff observed in this study would be a natural extension (Lewis and others, 2020).
Third, evaluating larger open-source and proprietary models — when hardware resources permit — would clarify whether the independence between consistency and accuracy observed here persists at greater parameter scales, or whether it is specific to sub-7B models.
@@
Translationally, our findings support the development of a lightweight clinical decision support prototype for general practitioners in low-resource settings, built around the model demonstrating the strongest balance of accuracy and instruction adherence, with prompt engineering guardrails informed by our sensitivity analysis. Finally, extending this evaluation framework to multilingual datasets and non-English clinical benchmarks is critical, as many low-resource healthcare environments operate outside English-dominant settings. Together, these directions would advance this work from controlled benchmarking toward safe, deployable clinical AI systems grounded in both technical rigour and real-world constraints.
@@ 6 Conclusion
In this study, we systematically evaluated prompt sensitivity in five open-source language models (2B–7B parameters), including four general instruction-tuned models and one domain-adapted pretraining model, across three clinical question answering datasets using five controlled prompt style variations, yielding 3,000 total evaluation instances.
We report four principal findings. First, consistency and accuracy are independent metrics: models that achieve high consistency are not necessarily more accurate, with Gemma 2 demonstrating the most pronounced case of reliable incorrectness — achieving the highest consistency scores while recording the lowest accuracy across all datasets. Second, roleplay-style prompts consistently reduced accuracy across all models and datasets, suggesting that persona-based framing introduces systematic instability in clinical reasoning tasks and should be avoided in healthcare AI prompt design. Third, instruction-following failures varied substantially across models and were not determined by parameter count alone, indicating that model size is an insufficient proxy for deployment reliability under hardware constraints. Fourth, Meditron-7B, a domain-adapted model without instruction tuning, exhibited near-complete failure on structured clinical QA tasks despite encoding substantial medical domain knowledge — demonstrating that domain knowledge and instruction-following capability are orthogonal requirements for clinical AI deployment.
Taken together, these findings demonstrate that consistency alone cannot serve as a proxy for clinical reliability. Models that appear stable under prompt variation may still pose significant patient safety risks if their answers are systematically incorrect. Safe deployment of AI in clinical settings — particularly in low-resource environments where human oversight may be limited — requires joint evaluation of accuracy, consistency, and instruction adherence.
We call on the research community to move beyond accuracy-only evaluation frameworks for clinical AI and adopt multidimensional assessment approaches that reflect the reliability requirements of real-world healthcare deployment. The evaluation framework, datasets, and code introduced in this study are publicly available at https://github.com/shravani-01/clinical-llm-eval to support reproducibility and future research in this direction.

### v2

@@ 4.2 Effect of Prompt Style on Accuracy
Prompt style demonstrated a consistent and measurable effect on model accuracy across all five models and three datasets (Figure 3). The Roleplay prompt style, in which models were instructed to respond as a practicing physician, consistently underperformed relative to all other styles. The most pronounced decline was observed in Phi-3 Mini, which achieved 48.0% accuracy under the Direct style but only 26.5% under the Roleplay style on MedQA, representing a drop of 21.5 percentage points. Similar reductions were observed across Llama 3.2, Gemma 2, and Mistral 7B, suggesting that roleplay prompting is systematically detrimental rather than model-specific (Figure 4).
In contrast, the Direct and Original prompt styles produced the highest and most stable accuracy across models and datasets, suggesting that minimal framing is preferable for reliable clinical question answering. These findings carry direct implications for healthcare AI deployment: prompt formulations intended to simulate clinical expertise, such as persona-based instructions, may paradoxically reduce the reliability of small language model outputs in patient-facing applications.
@@ 4.3 Instruction-Following Failure Rate
The UNKNOWN rate, defined as the proportion of responses that did not contain a valid answer option, varied substantially across models and datasets (Figure 5). Phi-3 Mini exhibited the highest UNKNOWN rate on MedQA at 10.5%, indicating that approximately one in ten queries failed to produce a usable response. Mistral 7B, despite being the largest model evaluated, also demonstrated non-negligible UNKNOWN rates across all three datasets (4.7%–7.2%), suggesting that model size alone does not guarantee reliable instruction following.
In contrast, Llama 3.2 and Gemma 2 achieved the lowest UNKNOWN rates across datasets (0.8%–2.8% and 0.9%–2.1%, respectively), consistently producing responses in the expected format. These results suggest that instruction adherence is a model-specific characteristic independent of parameter count.
Meditron-7B exhibited the most severe instruction-following failures of all models evaluated, with UNKNOWN rates of 22.8% on MedQA, 22.4% on MedMCQA, and 99.0% on PubMedQA. The near-complete failure on PubMedQA, where 99 of 100 responses were invalid, is consistent with Meditron’s architecture as a base pretraining model without instruction tuning. Unlike the instruction-following failures observed in Phi-3 Mini, which appear to be prompt-style dependent, Meditron’s failures reflect a fundamental absence of structured response capability rather than prompt sensitivity.
@@ 4.4 Consistency Versus Accuracy Relationship
Analysis of the relationship between mean consistency scores and overall accuracy revealed no clear positive correlation across models and datasets (Figure 6). Gemma 2 exemplifies this disconnect most clearly: it achieved the highest consistency scores across all datasets (0.845–0.888) while simultaneously recording the lowest accuracy (33.0–43.5%), indicating that it consistently produces incorrect answers with high confidence. Conversely, Llama 3.2 demonstrated moderate consistency (0.774–0.807) yet achieved the highest accuracy across most datasets (49.0–65.0%), showing that a model can be more frequently correct even when its answers vary across prompt styles.
The distribution of consistency scores further illustrates model-level differences in reliability (Figure 7). Phi-3 Mini exhibited the widest spread of consistency scores on MedQA, indicating substantial variability in prompt sensitivity across questions. Gemma 2 showed a tighter, higher distribution, remaining consistently stable but systematically inaccurate.
@@ 5.1 Consistency and Accuracy are Independent Metrics
A central finding of this study is that consistency and accuracy represent independent dimensions of small language model performance in clinical question answering. While prior evaluations of clinical AI have predominantly focused on accuracy as the primary metric (Bedi and others, 2025), the results presented here demonstrate that a model can exhibit high consistency while still producing systematically incorrect answers, a failure mode referred to in this work as reliable incorrectness.
Gemma 2 illustrates this most clearly, achieving the highest mean consistency scores across all datasets (0.845–0.888) while recording the lowest overall accuracy (33.0–43.5%). This pattern suggests that the model has learned stable response tendencies that are nonetheless clinically incorrect.
The danger of reliable incorrectness is both psychological and technical. Consider a primary care physician using an AI assistant to support diagnostic reasoning. If the model returns the same answer regardless of how a question is phrased, this repetition creates an illusion of reliability. A clinician might reason, “it keeps suggesting pulmonary embolism; it must be confident.” That apparent stability can feel reassuring. However, if the correct diagnosis is pneumonia, the model’s consistency actively reinforces a false conclusion. Over time, clinicians may develop misplaced trust in systems that appear stable and decisive without recognising that the stability reflects systematic error rather than clinical knowledge. In practice, this could translate into repeated misdiagnoses, inappropriate investigations, or delayed treatment, not because the model behaves randomly, but because it is confidently and predictably incorrect.
This finding aligns with Kim et al. (Kim and others, 2025), who demonstrate that foundation models frequently produce hallucinations with high apparent confidence and that clinicians consider such outputs particularly dangerous because they are difficult to detect. The present results extend this concern to the domain of prompt sensitivity. A model that produces consistent but incorrect answers across varied prompt formulations compounds this risk by appearing robust to rephrasing.
These observations underscore that accuracy and consistency must be evaluated jointly in any clinical AI assessment framework. Neither metric alone is sufficient to characterize the safety profile of a model intended for healthcare deployment.
@@ 5.2 Roleplay Prompts Consistently Underperform
Roleplay prompts appear to introduce a form of task interference that is particularly detrimental for small language models. When a prompt begins with persona framing such as “You are a senior physician taking a licensing examination,” the model must simultaneously interpret and simulate a clinical identity while reasoning through a structured medical question.
For smaller models with limited representational capacity, this dual requirement may dilute attention from the core reasoning task. Furthermore, persona-based prompts may activate patterns learned from conversational or narrative training data rather than the structured exam-style reasoning required for clinical QA benchmarks, effectively shifting the task from question answering to performance-oriented text generation.
This finding has direct practical implications. Prompt engineering guidelines for clinical AI systems often recommend persona framing to make outputs sound more authoritative or clinically appropriate (White and others, 2023). These results suggest that, for small open-source models, such framing may paradoxically reduce factual reliability. Developers building clinical decision support tools on small language models should favor minimal, direct prompt formulations over persona-based instructions.
This observation also extends findings from Ngweta et al. (Ngweta and others, 2024), who demonstrated that prompt format changes cause measurable performance variation in general NLP tasks. The present results show that, in clinical settings, one specific prompt format, roleplay, is systematically and substantially more harmful than others, representing a concrete and actionable finding for healthcare AI deployment.
@@ 5.3 Implications for Low-Resource Healthcare Deployment
A primary motivation for this study was the practical challenge of deploying AI in resource-constrained healthcare environments. Many rural clinics, community hospitals, and health systems in low-income settings lack access to GPU infrastructure, reliable high-speed internet, or cloud-based AI services. In these contexts, locally deployable open-source models in the 2B–7B parameter range represent a more realistic and equitable alternative to large proprietary systems. The experimental design intentionally reflected these constraints: all models were evaluated on consumer CPU hardware without domain-specific fine-tuning, simulating realistic deployment conditions rather than idealized research environments.
These findings provide concrete guidance for model selection in such settings. Although Gemma 2 demonstrated the highest consistency scores, its substantially lower accuracy suggests that it may systematically mislead clinicians, particularly in low-oversight environments where AI outputs may not be routinely verified. Mistral 7B, despite being the largest model evaluated, did not clearly outperform smaller alternatives and exhibited unexpected instruction-following failures, suggesting that larger parameter counts do not automatically justify greater hardware demands.
Llama 3.2 (3B) emerged as the most balanced candidate under the evaluation criteria used in this study, achieving the highest overall accuracy (49.0–65.0%), moderate consistency, and among the lowest instruction-following failure rates. For low-resource healthcare deployment scenarios where clinician oversight may be limited, a model offering strong accuracy and reliable instruction adherence is preferable to one that maximizes consistency at the expense of correctness.
More broadly, these results suggest that deployment decisions for clinical AI systems should be guided by joint evaluation of accuracy, consistency, and instruction compliance rather than any single metric. In settings where errors carry direct patient safety implications, the cost of systematic incorrectness, even when delivered consistently, outweighs the benefit of apparent stability. This study provides a practical evaluation framework for low-resource healthcare AI deployment that can be reproduced without specialized infrastructure, making it accessible to researchers and practitioners in the settings it is intended to serve.
@@ 5.4 Domain Knowledge Versus Instruction Following
This finding suggests that domain knowledge and task usability are orthogonal capabilities. A model can encode rich medical knowledge through continued pretraining yet remain clinically unusable for structured question answering if it lacks the ability to follow task-specific instructions. For practitioners deploying AI in clinical settings, this distinction has direct practical implications. Selecting a model based solely on its training data, without verifying instruction-following capability, may result in a system that is knowledgeable but functionally unusable in structured workflows.
These observations also motivate a clear direction for future work: instruction-tuning Meditron or similar domain-pretrained models on structured clinical QA datasets and evaluating whether the combination of domain knowledge and instruction following produces models that are both more accurate and more consistent than general instruction-tuned alternatives.
@@ 5.5 Limitations
Several limitations should be considered when interpreting these findings. First, 200 questions per dataset were sampled, which provides a computationally feasible and reproducible evaluation but may not fully capture the diversity of each benchmark. Second, the three datasets used, MedQA, MedMCQA, and PubMedQA, focus primarily on structured examination-style and research-based questions and may not reflect the complexity of real-world clinical dialogue or unstructured patient interactions.
Third, human clinical evaluation was not conducted to assess whether consistent but incorrect model outputs would meaningfully influence physician decision-making in practice. Such evaluation would strengthen the clinical validity of these findings but was beyond the scope of this study. Fourth, all models were evaluated without domain-specific fine-tuning, which may underestimate the performance achievable through medical adaptation techniques such as instruction tuning or retrieval-augmented generation (Lewis and others, 2020).
Fifth, the evaluation was restricted to multiple-choice and yes/no/maybe answer formats, whereas real clinical environments frequently involve open-ended reasoning, differential diagnosis, and contextual patient data. Sixth, hardware constraints limited the evaluation to models in the 2B–7B parameter range running on consumer CPU hardware, precluding direct comparison with larger proprietary systems such as GPT-4 or Claude. Seventh, while the five prompt styles were designed to simulate realistic variation in clinical query formulation, they do not exhaust the full range of phrasings a clinician might use in practice. Finally, Meditron-7B was evaluated in its base pretraining form without instruction tuning, which limits direct comparison with instruction-tuned models and may underrepresent the potential of domain-adapted models when properly instruction-tuned for structured clinical tasks.
Together, these limitations suggest that the results should be interpreted as an initial systematic analysis of prompt sensitivity under constrained deployment conditions rather than a comprehensive evaluation of all clinical AI scenarios. Future work should address these gaps through larger samples, broader dataset coverage, and human clinical validation studies.
@@ 5.6 Future Work
Several directions emerge naturally from this study. First, future research should evaluate whether domain-specific fine-tuning of small models on medical corpora reduces prompt sensitivity and improves consistency. If fine-tuned models demonstrate stronger stability under prompt variation, this would suggest that training data alignment, rather than model architecture, is the primary driver of inconsistency.
In particular, instruction-tuning Meditron-7B or similar domain-pretrained models on structured clinical QA datasets represents a direct next step. This would test whether combining domain knowledge with instruction-following capability produces models that are both more accurate and more consistent than the general instruction-tuned alternatives evaluated in this study.
Second, integrating retrieval-augmented generation (RAG) pipelines with small clinical models represents a promising direction. RAG-augmented systems ground responses in retrieved evidence, which may reduce both factual hallucinations and prompt-induced answer variation. Evaluating whether external knowledge retrieval mitigates the consistency–accuracy tradeoff observed in this study would be a natural extension (Lewis and others, 2020).
Third, evaluating larger open-source and proprietary models, when hardware resources permit, would clarify whether the independence between consistency and accuracy observed here persists at greater parameter scales or whether it is specific to sub-7B models.
@@
Translationally, these findings support the development of a lightweight clinical decision support prototype for general practitioners in low-resource settings, built around the model demonstrating the strongest balance of accuracy and instruction adherence, with prompt engineering guardrails informed by the sensitivity analysis. Extending this evaluation framework to multilingual datasets and non-English clinical benchmarks is also critical, as many low-resource healthcare environments operate outside English-dominant settings. Together, these directions would advance this work from controlled benchmarking toward safe, deployable clinical AI systems grounded in both technical rigor and real-world constraints.
@@ 6 Conclusion
This study systematically evaluates prompt sensitivity in five open-source language models (2B–7B parameters), including four general instruction-tuned models and one domain-adapted pretraining model, across three clinical question answering datasets using five controlled prompt style variations, yielding 3,000 total evaluation instances.
Four principal findings emerge from this analysis. First, consistency and accuracy are independent metrics: models that achieve high consistency are not necessarily more accurate, with Gemma 2 demonstrating the most pronounced case of reliable incorrectness, achieving the highest consistency scores while recording the lowest accuracy across all datasets. Second, roleplay-style prompts consistently reduce accuracy across all models and datasets, suggesting that persona-based framing introduces systematic instability in clinical reasoning tasks and should be avoided in healthcare AI prompt design. Third, instruction-following failures vary substantially across models and are not determined by parameter count alone, indicating that model size is an insufficient proxy for deployment reliability under hardware constraints. Fourth, Meditron-7B, a domain-adapted model without instruction tuning, exhibited near-complete failure on structured clinical QA tasks despite encoding substantial medical domain knowledge, demonstrating that domain knowledge and instruction-following capability are orthogonal requirements for clinical AI deployment.
Taken together, these results demonstrate that consistency alone cannot serve as a proxy for clinical reliability. Models that appear stable under prompt variation may still pose significant patient safety risks if their answers are systematically incorrect. Safe deployment of AI in clinical settings, particularly in low-resource environments where human oversight may be limited, requires joint evaluation of accuracy, consistency, and instruction adherence.
More broadly, this work highlights the need for the research community to move beyond accuracy-only evaluation frameworks for clinical AI and adopt multidimensional assessment approaches that reflect the reliability requirements of real-world healthcare deployment. The evaluation framework, datasets, and code introduced in this study are publicly available at (Hariprasad, 2026) to support reproducibility and future research in this direction.